# On $\operatorname{CAT}(\kappa)$ surfaces

### v2

@@ 3.1. Basic definitions
defined for all $(x,y)\in\mathbb{R}^{2}$, with the additional requirement that $\sqrt{x^{2}+y^{2}}<1/|\kappa|$ if $\kappa<0$. If $\kappa=0$, this formula gives a rescaling of the usual Euclidean metric, but the resulting metric space is of course isometric to the Euclidean plane. If $\kappa>0$, then we must add in the point $\infty$ to obtain the complete $2$-sphere. One immediate consequence is that any two model spaces are locally bi-Lipschitz equivalent. More precisely, let $\kappa_{1},\kappa_{2}\in\mathbb{R}$, $x_{1}\in M_{\kappa_{1}}^{2}$ and $x_{2}\in M_{\kappa_{2}}^{2}$, and $0<r<\min\{D_{\kappa_{1}},D_{\kappa_{2}}\}$. Then the balls $B(x_{1},r)$ and $B(x_{2},r)$ are $L$-bi-Lipschitz equivalent, where $L\to 1$ uniformly (after fixing $\kappa_{1},\kappa_{2}$) as $r\to 0$. In particular, for all $x_{1}\in M_{0}^{2}$ and $x_{2}\in M_{1}^{2}$, the ball $B(x_{1},2)\subset M_{0}^{2}$ is $2$-bi-Lipschitz equivalent to the ball $B(x_{2},\pi/2)\subset M_{1}^{2}$.
@@
Given a point $x\in[pq]$, we let $\bar{x}$ denote the point on $[\bar{p}\bar{q}]$ such that $d(p,x)=\bar{d}_{\kappa}(\bar{p},\bar{x})$ and $d(x,q)=\bar{d}_{\kappa}(\bar{x},\bar{q})$. A triangle $T=\triangle pqr\subset X$ with perimeter at most $2D_{\kappa}$ is said to satisfy the $\operatorname{CAT}(\kappa$) inequality if
@@ 3.2. Angles
Let $X$ be an arbitrary metric space. Given a triangle $T=\triangle pqr\subset X$, the model angle at $p$, denoted ¯$\bar{\angle}_{p}^{\kappa}(q,r)$, is the angle between the geodesics $[\bar{p}\bar{q}]$ and $[\bar{p}\bar{r}]$ in the comparison triangle $\triangle\bar{p}\bar{q}\bar{r}$ in $M^{2}_{\kappa}$. This is defined provided that the perimeter of $T$ is at most $2D_{\kappa}$. In a $\operatorname{CAT}(\kappa)$ space, the model angle satisfies the triangle inequality: ¯$\bar{\angle}_{p}^{\kappa}(q,r)\leq\bar{\angle}_{p}^{\kappa}(q,s)+\bar{\angle}_{p}^{\kappa}(s,r)$ for all $p,q,r,s$, provided that all these angles are defined.
@@
[…]
Let $T$ be a triangle with angles $\alpha,\beta,\gamma$. The angle excess or excess of $T$ is $\delta(T)=\alpha+\beta+\gamma-\pi$. If $\mathcal{T}$ is a finite family of geodesic triangles, the excess of $\mathcal{T}$ is $\delta(\mathcal{T})=\sum_{i}\delta(T_{i})$. In connection with excess, we recall the classical Girard’s theorem, which states that the area of a triangle in the $2$-sphere $M_{1}^{2}$ is precisely equal to its excess. More generally, if $T$ is a triangle in the model space $M_{\kappa}^{2}$ for $\kappa>0$, then $|T|=\delta(T)/\kappa$.
3.3. Spherical law of cosines
In this section, we restrict to the case where $\kappa>0$. We recall the spherical law of cosines. See for example Section 1.A of [2]. Consider a triangle with side lengths $a,b,c$ and corresponding opposite angles $\alpha,\beta,\gamma$ in the model space $M_{\kappa}^{2}$. Then
$\cos\gamma=\frac{\cos(c\sqrt{\kappa})-\cos(a\sqrt{\kappa})\cos(b\sqrt{\kappa})}{\sin(a\sqrt{\kappa})\sin(b\sqrt{\kappa})}.$
By expanding this equation with power series, we obtain the following inequality for fixed $\kappa>0$: for all $\varepsilon>0$, there exists $\delta>0$ such that if $a,b<\delta$ then
(3.2) $(1-\varepsilon)(a^{2}+b^{2}-2ab\cos\gamma)\leq c^{2}\leq(1+\varepsilon)(a^{2}+b^{2}-2ab\cos\gamma).$
As a particular consequence, if $\gamma\geq\theta$ for some fixed value $\theta>0$, then there is a corresponding constant $C_{\theta}>0$ such that
(3.3) $C_{\theta}\min\{a^{2},b^{2}\}\leq(1-\varepsilon)(a^{2}+b^{2}-2ab\cos\gamma)\leq c^{2}$
whenever $a,b<\delta$.
The following lemma forms the basis of the second proof of 1.1.
@@ 3.3. Spherical law of cosines
Fix $\kappa>0$. For all $\theta>0$, there exists $L\geq 1,\delta>0$ such that the following holds for every triangle $T$ in a $\operatorname{CAT}(\kappa)$ space with each angle at least $\theta$ and having perimeter at most $\delta$.
  1. (1)
    The ratio of the length of any two edges of $T$ is bounded by $L$.
  2. (2)
    The canonical map from $T$ to its Euclidean model triangle $\bar{T}_{0}$ is $L$-bi-Lipschitz.
  3. (3)
    If $T$ is a Jordan triangle, then
    $\delta(T)\leq 4|\bar{T}_{0}|\leq 16L^{2}|T|.$
The canonical map in (2) is the map that sends each edge of $T$ to the corresponding edge of $\bar{T}_{0}$ preserving arc length, that is, each point $x\in T$ to the point $\bar{x}\in\bar{T}_{0}$.
@@
By rescaling, we may assume that $\kappa=1$. Choose $\delta>0$ small enough so that (3.2) holds for $\varepsilon=1/2$ for any triangle in $M_{\kappa}^{2}$ with perimeter at most $\delta$. Assume further that $\delta\leq\pi/2$. Consider a triangle $T$ with vertices $p,q,r$, with each angle at least $\theta$ and perimeter at most $\delta$. Statement (1) follows immediately from (3.3).
We now verify statement (2). Let $\bar{T}_{1}$ be the spherical model triangle corresponding to $T$ and let $\bar{T}_{0}$ be the Euclidean model triangle. Let $x,y\in T$. If $x,y$ belong to the same side, then $d(x,y)=\bar{d}_{1}(\bar{x},\bar{y})$ and there is nothing more to prove. Otherwise, by symmetry, we may assume that the sides containing $x$ and $y$, respectively, meet at a common vertex $p$. Let $\theta_{0}$ denote the angle at the vertex $\bar{p}$ in $\bar{T}_{1}$. Let $\bar{T}=\triangle\bar{p}^{\prime}\bar{x}^{\prime}\bar{y}^{\prime}$ denote the spherical model triangle corresponding to the triangle $\triangle pxy$, and $\bar{\theta}$ the angle at $\bar{p}^{\prime}$. The $\operatorname{CAT}(\kappa)$ condition implies
$d(x,y)=\bar{d}_{1}(\bar{x}^{\prime},\bar{y}^{\prime})\leq\bar{d}_{1}(\bar{x},\bar{y})$
and moreover that $\theta_{0}\geq\bar{\theta}\geq\theta$. For the following inequalities, we use the auxiliary function
$L(\rho)=d(p,x)^{2}+d(p,y)^{2}-2d(p,x)d(p,y)\cos(\rho).$
Applying (3.2) with the value $\varepsilon=1/2$ gives
$(1/2)L(\bar{\theta})\leq\bar{d}_{1}(\bar{x}^{\prime},\bar{y}^{\prime})^{2}=d(x,y)^{2}.$
Since $\theta\leq\bar{\theta}$, the left-hand side of the above inequality is bounded below by $(1/2)L(\theta)$, which in turn has a lower bound of the form $C_{\theta}^{\prime}(d(p,x)^{2}+d(p,y)^{2})$, $C_{\theta}^{\prime}$ a constant depending only on $\theta$. On the other hand, we have the upper bound $L(\theta_{0})\leq d(p,x)^{2}+d(p,y)^{2}$. By (3.2), we have $L(\theta_{0})\geq(2/3)\bar{d}_{1}(\bar{x},\bar{y})^{2}$. Combining these inequalities gives
$\frac{L(\bar{\theta})}{2}=\frac{L(\bar{\theta})}{2L(\theta_{0})}L(\theta_{0})\geq\frac{C_{\theta}^{\prime}}{2}\cdot\frac{2}{3}\bar{d}_{1}(\bar{x},\bar{y})^{2}.$
We conclude that $\bar{d}_{1}(\bar{x},\bar{y})\leq\sqrt{3/C_{\theta}^{\prime}}\cdot d(x,y)$, hence that the canonical map from $T$ to $\bar{T}_{1}$ is bi-Lipschitz with constant only depending on $\theta$. Observe that, since the perimeter of $T$ is at most $\pi/2$, the triangles $\bar{T}_{0}$ and $\bar{T}_{1}$ are $2$-bi-Lipschitz equivalent. Then (2) follows immediately.
Finally, we show that (3) holds as a consequence of the Besicovitch inequality (2.1). Recall that the triangles $\bar{T}_{0}$ and $\bar{T}_{1}$ are $2$-bi-Lipschitz equivalent. Applying the Besicovitch inequality and Girard’s theorem, we see that
$|T|\geq\frac{|\bar{T}_{0}|}{4L^{2}}\geq\frac{|\bar{T}_{1}|}{16L^{2}}=\frac{\delta(\bar{T}_{1})}{16L^{2}}\geq\frac{\delta(T)}{16L^{2}}.$
This verifies (3). ∎
A similar statement, with more intricate proof, holds for surfaces of bounded curvature. See Lemma 4 in [4].
3.4. Facts about $\operatorname{CAT}(\kappa)$ spaces
We record a few foundational facts about $\operatorname{CAT}(\kappa)$ spaces. The first of these is Reshetnyak’s majorization theorem; see Section 9.L of [2].
Let $D\subset M_{\kappa}^{2}$ be a closed Jordan domain whose boundary has finite length, and let $X$ be a metric space. A map $f\colon D\to X$ is majorizing if it is $1$-Lipschitz and its restriction to the boundary $\partial D$ is length-preserving. If $\Gamma$ is a closed curve in $X$, we say that $D$ majorizes $\Gamma$ if there is a majorizing map $f\colon D\to X$ such that the restriction $f|_{\partial D}$ traces out $\Gamma$.
Theorem 3.2 (Reshetnyak’s majorization theorem).
For any closed curve $\Gamma$ in a $\operatorname{CAT}(\kappa)$ space of length at most $2D_{\kappa}$, there exists a convex region $D$ in $M^{2}_{\kappa}$, and an associated map $f$ such that $D$ majorizes $\Gamma$ under $f$.
If $\Gamma$ itself is the boundary of a simple triangle $T$, then it is easy to see that $D$ is the model triangle $\bar{T}_{\kappa}$.
A consequence of Reshetnyak’s majorization theorem is that any $\operatorname{CAT}(\kappa)$ surface has locally finite Hausdorff $2$-measure. Consider a triangulation $\mathcal{T}$ of the surface as given by 2.5, where $\varepsilon$ is chosen to satisfy $\varepsilon<D_{\kappa}/2$ and each triangle is contained in a $\operatorname{CAT}(\kappa)$ neighborhood. For each triangle $T\in\mathcal{T}$, Reshetnyak’s theorem gives a $1$-Lipschitz map $\bar{T}_{1}\to T$. Since $\bar{T}_{1}$ has finite area and $\mathcal{T}$ is locally finite, we see that the Hausdorff $2$-measure on $X$ is locally finite. Note however that we do not need the full strength of the triangulation theorem for this conclusion, but rather just the much easier property that any point lies in a polygonal neighborhood (Lemma 5.1 in [16]).
The next fact about $\operatorname{CAT}(\kappa)$ spaces concerns uniqueness of geodesics. If two points $x,y\in X$ satisfying $d(x,y)<D_{\kappa}$ are joined by a geodesic, then that geodesic is unique. The existence of a geodesic between sufficiently close points is a consequence of the Hopf–Rinow theorem; see Proposition 3.7 in [9]. Suppose that $x,y$ belong to a neighborhood $U\subset X$ homeomorphic to a disk. If $d(x,y)\leq d(x,\partial U)$, then there is a geodesic in $X$ from $x$ to $y$. Moreover, the ball $B(x,r)$ is convex and contractible for all $r<\min\{d(x,\partial U)/2,D_{\kappa}/2\}$.
@@ 3.4. Facts about $\operatorname{CAT}(\kappa)$ spaces
Let $X$ be a $\operatorname{CAT}(\kappa)$ space that is also a topological manifold. Then every geodesic can be extended indefinitely in both directions as a local geodesic. It is length-minimizing for all pairs of points at distance less than $D_{\kappa}$.
See Proposition II.5.125.125.125.12 in [9] for a proof.
Corollary 3.4.
Let $T=\triangle pqr$ be a Jordan triangle and $x$ an interior point. Then the geodesic $[px]$ can be extended to intersect the opposite side $[qr]$.
@@
By 3.3, $[px]$ can be extended indefinitely beyond $x$. Since the triangle $T$ is compact, this extension must intersect the boundary $\partial T$ at some point. This point must lie on the side $[yz]$ since otherwise the uniqueness of geodesic is violated. ∎
As a consequence of geodesic extendability, we can show that small triangles in $\operatorname{CAT}(\kappa)$ spaces are convex.
Lemma 3.5.
Let $T$ be a Jordan triangle in a $\operatorname{CAT}(\kappa)$ space with perimeter at most $D_{\kappa}/2$. Assume further that $T$ is contained in a neighborhood $U$ homeomorphic to an open disk with $d(T,\partial U)\geq\ell(\partial T)$. Then $T$ is convex.
Proof.
Consider two points $x,y\in T$. Let $\Gamma$ be the geodesic between them, which exists by the second assumption. We must show that the image of $\Gamma$ is contained in $T$. If $x,y$ belong to the same side of $T$, then $\Gamma$ is just a subarc of this side and we are done. Assume then that $x$ belongs to the side $A$ and $y$ belongs to the side $B$, with $C$ the other side of $T$. Extend the side $A$ in both directions until reaching $\partial U$. This divides $U$ into two non-overlapping closed disks $U_{A}^{1}$ and $U_{A}^{2}$, where $T\subset U_{A}^{1}$. Do the same for $B$ to obtain two non-overlapping closed disks $U_{B}^{1}$ and $U_{B}^{2}$. If $\Gamma$ leaves the set $U_{A}^{1}$, then we homotope any subcurves of $\Gamma$ in $(U_{A}^{2})^{\circ}$ onto $A$ to obtain a curve $\Gamma_{1}$ in $U_{A}^{1}$. It is clear that $\ell(\Gamma_{1})\leq\ell(\Gamma)$. Next, if $\Gamma_{1}$ leaves the set $U_{B}^{1}$, we do the same procedure to obtain a curve $\Gamma_{2}$ in $U_{B}^{1}\cap U_{A}^{1}$. This curve necessarily satisfies $\ell(\Gamma_{2})\leq\ell(\Gamma_{1})$. Finally, if the curve $\Gamma_{2}$ leaves $T$, it must be through the side $C$. We homotope any subcurve of $\Gamma_{2}$ in $U\setminus T$ onto $C$ to obtain a curve $\Gamma_{3}$ contained in $T$. It is clear that $\ell(\Gamma_{3})\leq\ell(\Gamma_{2})$. ∎
@@ 3.5. Surfaces of bounded curvature
Surfaces of bounded curvature can be defined equivalently in terms of a sequence of polyhedral approximations. Namely, a length surface $(X,d)$ has bounded curvature if and only if there is a sequence of polyhedral metrics $d_{n}$ on $X$ with locally uniformly bounded curvature that converges uniformly to $d$. The same statement is true with “polyhedral metric” replaced by “Riemannian metric”; in this case, the curvature measure is defined by integrating the Gaussian curvature with respect to area. Surfaces of bounded curvature can also be characterized in terms of the existence of conformal parametrizations due to Reshetnyak, which we discussed briefly in Section 1.3.
The characterization as limits of polyhedral surfaces enables one to define area and the curvature measure on a surface of bounded curvature by taking weak limits. We call the notion of area obtained this way the Alexandrov area and denote the area of a set $A$ by $\operatorname{Area}(A)$. More precisely, let $(d_{n})$ be a sequence of polyhedral metrics on $X$ with locally uniformly bounded curvature converging uniformly to a metric $d$, with $\varphi_{n}$ denoting the transition map from $(X,d_{n})$ to $(X,d)$ and $\omega_{n}$ the curvature measure for $d_{n}$. For all $A\subset X$ we define $\operatorname{Area}(A)=\lim_{n\to\infty}|\varphi_{n}^{-1}(A)|$. Likewise, we define $\omega(A)=\lim_{n\to\infty}\omega_{n}(\varphi_{n}^{-1}(A))$ It can be shown that these are independent of the choice of approximating surface and hence well-defined; see Section VIII.3 of [1].
We mention a few more advanced concepts related to surfaces of bounded curvature. First is the notion of sector angle. Let $p\in X$ be a point, let $[pq]$ and $[pr]$ be geodesics intersecting only at $p$, and let $O$ be a closed oriented Jordan neighborhood of $p$ divided into two sectors $U$ and $V$. One of these, say $U$, contains the arc of $\partial O$ running counterclockwise from $[pq]$ to $[pr]$. We define the sector angle of $[pq]$ and $[pr]$ as follows. Let $a_{0},\ldots,a_{n}$ be a sequence of points in $\partial U\cap\partial O$ in cyclic order with $a_{0}\in[pq]$ and $a_{n}\in[pr]$, and $[pa_{1}],\ldots,[pa_{n}]$ be a sequence of geodesics intersecting $\partial U$ only at their endpoints and intersecting each other only at $p$. The sector angle of $[pq]$ and $[pr]$, denoted by ~$\widetilde{\angle}_{p}(q,r)$, is the supremum of the sum $\sum_{i=1}^{n}\angle_{p}(a_{i},a_{i+1})$, where the supremum is taken over all such $n$-tuples $a_{0},\ldots,a_{n}$ and admissible choices of geodesics. The sector angle is independent of the choice of neighborhood $O$.
Next, the total angle at a point p is the supremum of the sum of sector angles taken over all partitions of a neighborhood of $p$ into sectors. We record the observation that if $p$ is the interior point of a geodesic, then the sector angle at $p$ is at least $2\pi$, since the sector corresponding to each side of the geodesic contributes at least $\pi$ to the total angle.
Let $X$ be a surface of bounded curvature and let $D$ be a Jordan polygon. Denote the sector angle at a vertex $v_{i}$ of $\partial D$ on the interior side of $\partial D$ by $\phi_{i}$. Assume further that the interior sector angle of each non-vertex point of $\partial D$ is $\pi$. We define the interior rotation of the curve $\partial D$ by $\tau(\partial D)=\sum_{i}(\pi-\phi_{i})$. The interior rotation can be defined for more general simple closed curves by taking limits of polygons, though we do not need this generality here.
The following version of the Gauss–Bonnet theorem holds for surfaces of bounded curvature. See Theorem VI.5 of [1].
Theorem 3.7 (Gauss–Bonnet).
If $X$ is a surface of bounded curvature and $D\subset X$ is a polygon, then
$\tau(\partial D)+\omega(D^{\circ})=2\pi.$

### v3

@@ 3.1. Basic definitions
defined for all $(x,y)\in\mathbb{R}^{2}$, with the additional requirement that $x^{2}+y^{2}<1/|\kappa|$ if $\kappa<0$. If $\kappa=0$, this formula gives a rescaling of the usual Euclidean metric, but the resulting metric space is of course isometric to the Euclidean plane. If $\kappa>0$, then we must add in the point $\infty$ to obtain the complete $2$-sphere.
@@
Given a point $x\in[pq]$, we let $\bar{x}$ denote the point on $\bar{p}\bar{q}]$ such that $d(p,x)=\bar{d}_{\kappa}(\bar{p},\bar{x})$ and $d(x,q)=\bar{d}_{\kappa}(\bar{x},\bar{q})$. A triangle $T=\triangle pqr\subset X$ with perimeter at most $2D_{\kappa}$ is said to satisfy the $\operatorname{CAT}(\kappa$) inequality if
@@ 3.2. Angles
Let $X$ be an arbitrary metric space. Given a triangle $T=\triangle pqr\subset X$, the model angle at $p$, denoted $\bar{\angle}_{p}^{\kappa}(q,r)$, is the angle between the geodesics $\bar{p}\bar{q}]$ and $\bar{p}\bar{r}]$ in the comparison triangle $\triangle\bar{p}\bar{q}\bar{r}$ in $M^{2}_{\kappa}$. This is defined provided that the perimeter of $T$ is at most $2D_{\kappa}$. In a $\operatorname{CAT}(\kappa)$ space, the model angle satisfies the triangle inequality: $\bar{\angle}_{p}^{\kappa}(q,r)\leq\bar{\angle}_{p}^{\kappa}(q,s)+\bar{\angle}_{p}^{\kappa}(s,r)$ for all $p,q,r,s$, provided that all these angles are defined.
@@
where $q^{\prime}\in[pq]\setminus\{p\}$ and $r^{\prime}\in[pr]\setminus\{p\}$. In general, the upper angle satisfies the triangle inequality: $\angle_{p}(q,r)\leq\angle_{p}(q,s)+\angle_{p}(s,r)$ for all points $p,q,r,s$ such that the geodesics $[pq],[pr],[ps]$ exist. If the $\limsup$” in (3.1) is an actual limit, then $\angle_{p}(q,r)$ is called the angle between $[pq]$ and $[pr]$. If $X$ is a $\operatorname{CAT}(\kappa)$ space, the angle between adjacent sides of any geodesic triangle of perimeter less than $2D_{\kappa}$ always exists and is less than or equal to the corresponding model angle.
Let $T$ be a triangle with angles $\alpha,\beta,\gamma$. The angle excess or excess of $T$ is $\delta(T)=\alpha+\beta+\gamma-\pi$. If $\mathcal{T}$ is a finite family of geodesic triangles, the excess of $\mathcal{T}$ is $\delta(\mathcal{T})=\sum_{i}\delta(T_{i})$. In connection with excess, we recall the classical Girard’s theorem, which states that the area of a triangle in the $2$-sphere $M_{1}^{2}$ is precisely equal to its excess. More generally, if $T$ is a triangle in the model space $M_{\kappa}^{2}$ for any $\kappa\in\mathbb{R}$, then $\kappa|T|=\delta(T)$.
3.3. Facts about $\operatorname{CAT}(\kappa)$ spaces
We record some properties of $\operatorname{CAT}(\kappa)$ spaces. The first concerns uniqueness of geodesics. If two points $x,y\in X$ satisfying $d(x,y)<D_{\kappa}$ are joined by a geodesic, then that geodesic is unique. The existence of a geodesic between sufficiently close points is a consequence of the Hopf–Rinow theorem; see Proposition I.3.7 in [7]. Suppose that $x,y$ belong to a neighborhood $U\subset X$ homeomorphic to a closed disk. If $d(x,y)\leq d(x,\partial U)$, then there is a geodesic in $X$ from $x$ to $y$. Moreover, the ball $B(x,r)$ is convex and contractible for all $r<\min\{d(x,\partial U)/2,D_{\kappa}/2\}$. See Proposition II.1.4 in [7].
@@ 3.3. Facts about $\operatorname{CAT}(\kappa)$ spaces
Let $X$ be a $\operatorname{CAT}(\kappa)$ space that is also a topological manifold. Then every geodesic can be extended indefinitely in both directions as a local geodesic. It is length-minimizing for all pairs of points at distance at most $D_{\kappa}$.
See Proposition II.5.125.12 in [7] for a proof. We can use geodesic extendability to show to following.
Corollary 3.2.
Let $X$ be a $\operatorname{CAT}(\kappa)$ space that is also a topological surface. Let $T=\triangle pqr$ be a Jordan triangle in $X$ of diameter less than $D_{\kappa}$ and let $x$ be an interior point. Then the geodesic $[px]$ can be extended to intersect the opposite side $[qr]$.
@@
By ˜3.1, $[px]$ can be extended indefinitely past the point $x$. Let $\Gamma$ denote a maximal extension past $x$, which we assume is parametrized by arc length. We claim that $\Gamma$ must leave the set $T$. Assume this is not the case. If $\Gamma$ has finite length, then its domain must be an interval of the form $[a,b)$. Since $\Gamma$ is $1$-Lipschitz and $T$ is compact, $\lim_{t\to b}\Gamma(t)$ must exist and lie in $T$. If we set $\Gamma(b)=\lim_{t\to b}\Gamma(t)$, then by continuity $\Gamma$ retains the property of being a geodesic. This contradicts the maximality of $\Gamma$. Hence $\Gamma$ must have infinite length. However, since $T$ has diameter at most $D_{\kappa}$, it must be the case that $\Gamma$ is length minimizing between all pairs of points it contains. Again since $T$ has diameter at most $D_{\kappa}$, this contradicts the statement that $\Gamma$ has infinite length. We conclude that $\Gamma$ intersects $\partial T$ at some point other than $p$. This point must lie on the side $[yz]$ since otherwise the uniqueness of geodesic is violated. ∎
Likewise, small Jordan triangles in $\operatorname{CAT}(\kappa)$ surfaces are convex:
@@
Let $X$ be a $\operatorname{CAT}(\kappa)$ space that is also a topological surface, and let $T$ be a Jordan triangle in $X$. Assume further that $T$ is contained in a neighborhood $U$ homeomorphic to a closed disk with $d(T,\partial U)\geq\ell(\partial T)$ and $\text{diam}(U)<D_{\kappa}$. Then $T$ is convex.
@@
Consider two points $x,y\in\partial T$. Observe that $d(x,y)<\ell(\partial T)$ and that $B(x,\ell(\partial T))\subset U$. This implies that there is a geodesic $\Gamma$ from $x$ to $y$ in $U$, which moreover is unique since $\operatorname{diam}(U)<D_{\kappa}$. We must show that $\Gamma$ is contained in $T$. If $x,y$ belong to the same side of $T$, then $\Gamma$ is just a subarc of this side and we are done. Assume then that $x$ belongs to the side $A$ and $y$ belongs to the side $B$, with $C$ the other side of $T$. By arguing as in ˜3.2, the geodesic $A$ can be extended in each direction until reaching the boundary $\partial U$. Since $\operatorname{diam}(U)<D_{\kappa}$, this extension is also a geodesic, which we still denote by $A$. This extension divides $U$ into two non-overlapping closed disks $U_{A}^{1}$ and $U_{A}^{2}$, where $T\subset U_{A}^{1}$. Do the same for $B$ to obtain two non-overlapping closed disks $U_{B}^{1}$ and $U_{B}^{2}$, with $T$ contained in $U_{B}^{1}$, and likewise for $C$. By the uniqueness of geodesics, $\Gamma$ cannot cross the geodesic $A$, and hence $\Gamma$ must be contained in the set $U_{A}^{1}$. Likewise, $\Gamma$ must be contained in the sets $U_{B}^{1}$ and $U_{C}^{1}$. Since $T=U_{A}^{1}\cap U_{B}^{1}\cap U_{C}^{1}$, we deduce that $\Gamma$ lies in $T$. Thus $T$ is convex. ∎
@@ 3.4. Surfaces of bounded curvature
Surfaces of bounded curvature can be defined equivalently in terms of a sequence of polyhedral approximations. Namely, a length surface $(X,d)$ has bounded curvature if and only if there is a sequence of polyhedral metrics $d_{n}$ on $X$ with locally uniformly bounded curvature that converges uniformly to $d$. The same statement is true with “polyhedral metric” replaced by “Riemannian metric”; in this case, the curvature measure is defined by integrating the Gaussian curvature with respect to area. Surfaces of bounded curvature can also be characterized in terms of the existence of conformal parametrizations due to Reshetnyak, which we discussed briefly in Section˜1.3.
The characterization as limits of polyhedral surfaces enables one to define area and the curvature measure on a surface of bounded curvature by taking weak limits. We call the notion of area obtained this way the Alexandrov area and denote the area of a set $A$ by $\operatorname{Area}(A)$. More precisely, let $(d_{n})$ be a sequence of polyhedral metrics on $X$ with locally uniformly bounded curvature converging uniformly to a metric $d$, with $\varphi_{n}$ denoting the transition map from $(X,d_{n})$ to $(X,d)$ and $\omega_{n}$ the curvature measure for $d_{n}$. For all Borel sets $A\subset X$ we define $\operatorname{Area}(A)=\lim_{n\to\infty}|\varphi_{n}^{-1}(A)|$. Likewise, we define $\omega(A)=\lim_{n\to\infty}\omega_{n}(\varphi_{n}^{-1}(A))$ It can be shown that these are independent of the choice of approximating surface and hence well-defined; see Section VIII.3 of [1].